OpenStudy (anonymous):

the sum of r consecutive positive integers will always be divisible by 2 if r is a multiple of 1. 6 2. 5 3. 4 4. 3 5. 2

OpenStudy (anonymous):

i dont know how to explain this one very well >.<

OpenStudy (anonymous):

Xand, do you know about looking at integers mod 2?

OpenStudy (anonymous):

If you dont, I guess one way to do this to try to think of counter examples for each answer. For example, if i thought 3 was the answer, i would pick: 2, 3, 4 as my counter example. That sum isnt divisible by 2, so it cant be the right answer. Just repeat this process until you have one answer left.

OpenStudy (zarkon):

can't we just look at \[(k)+(k+1)+(k+2)+\cdots (k+(r-1))\] \[=rk+\sum_{i=0}^{r-1}i\] \[=rk+\frac{(r-1)r}{2}\] then look at what happens when \[r=2a,r=3a,\ldots,r=6a\] we will be able to see that if \[r=4a\] then \[=rk+\frac{(r-1)r}{2}\] \[=4ak+\frac{(4a-1)4a}{2}=2\left(2ak+\frac{(4a-1)2a}{2}\right)\] \[=2\left(2ak+(4a-1)a\right)\]

OpenStudy (zarkon):

A real simple way of doing this: even + even is even odd +odd is even odd+even is odd so we want a sum of and even number of odd term clearly the only one that will guarantee this is a multiple of 4 since 2) o+e=o 3) e+o+e=o 4) e+o+e+o=e and o+e+o+e=e (both ways even) 5) o+e+o+e+o=o 6) o+e+o+e+o+e=o

OpenStudy (anonymous):

yep, that last post of yours is what i wanted to do with the mod 2 business. nice nice :)